OpenStudy (anonymous):

I think I have the answer-I just need to know if it's right? Write the product as a trinomial. (3r – 4)(r + 8)

OpenStudy (anonymous):

3r^2 – 28r – 32

OpenStudy (anonymous):

@amoodarya @beccaboo333 @tHe_FiZiCx99 @Compassionate

OpenStudy (beccaboo333):

I'm of no help

OpenStudy (amoodarya):

(3r – 4)(r + 8) = (3r )(r + 8) + ( – 4)(r + 8) = 3r^2+24r +(-4r -32) =3r^2+24r -4r-32 =3r^2+20r-32

OpenStudy (anonymous):

okay that makes sense, thank you so much!

OpenStudy (anonymous):

haha sorry @beccaboo333